# MSRS: Adaptive Multi-Subspace Representation Steering for Attribute Alignment in Large Language Models

### v3

@@ 6 Conclusion
We propose Multi-Subspace Representation Steering (MSRS), a framework for multi-attribute control in LLMs. MSRS combines attribute-specific subspaces, SVD-guided dimensionality, a shared subspace for cross-attribute correlations, and dynamic token selection to reduce steering conflicts and improve control over attributes like truthfulness, bias, instruction-following, refusal, and generation quality. Experiments demonstrate that MSRS outperforms existing methods, offering a scalable and robust solution for aligned language generation.
Appendix A LLM Usage Statement
In this research, Large Language Models (LLMs) were used to assist with tasks related to language refinement, formatting, and reference management. The LLM helped in improving the clarity and flow of the text, ensuring consistent citation formatting, and aiding in the organization of the manuscript. The authors affirm that all final content was reviewed and critically assessed by the authors to ensure the accuracy and relevance of the research.
Appendix B Implementation Details
A.1 Algorithm Figure 1(b) illustrates the overall structure of MSRS, which integrates shared and attribute-specific subspaces for precise and disentangled multi-attribute control. To complement this, we present the detailed training algorithm in Algorithm 1.
The MSRS training process begins by computing activation statistics from each attribute-specific dataset $\mathcal{D}_{i}$, aggregating token representations across examples to form mean activations $\tau_{i}$. These activations are used to extract a shared subspace $B_{\text{shared}}$ via singular value decomposition (SVD) over concatenated attribute activations, capturing the dominant, overlapping directions.
Next, we derive private subspaces $B_{i}$ for each attribute $i$ by removing the shared components from their individual activations, followed by a second SVD on the residuals. The final aligned subspace $S_{\text{align}}$ is constructed by concatenating the shared and private bases.
Training proceeds by optimizing the steerable representation function $\Phi_{l,p}(h;R,W,b,m)$, applied to token representations at a selected layer $l$ and position $p$. Once the intervention is applied, the resulting representation is passed through a Softmax function, which normalizes the output to produce a probability distribution over the possible classes. The Softmax function converts the raw scores (logits) produced by the intervention into a range between 0 and 1, ensuring that the sum of the outputs is equal to 1, representing the predicted class probabilities.The predicted probabilities are then compared to the ground truth labels $y$ using Cross-Entropy loss. Cross-Entropy measures the difference between the predicted probabilities (after applying Softmax) and the true label distribution. It quantifies the performance of the model by penalizing the difference, with the objective of minimizing this loss during training. Finally, the computed cross-entropy loss is assigned as the task loss $\mathcal{L}_{\text{task}}$, which is minimized through the optimization process. This ensures that the steerable representation function $\Phi_{l,p}(h;R,W,b,m)$ is optimized to produce more accurate predictions in subsequent tasks.
The total loss function $\mathcal{L}$ combines the task loss $\mathcal{L}_{\text{task}}$ with two regularization terms. The first term regularizes the mask network $m(h)$, encouraging it to stay close to a prior mask $m_{\text{prior}}$ through the $\ell_{2}$-norm term $\lambda_{1}\|m(h)-m_{\text{prior}}\|_{2}^{2}$. The second term enforces that the learned representation $R$ aligns with a reference matrix $S_{\text{align}}$ by minimizing the cosine dissimilarity, controlled by $\lambda_{2}\mathcal{L}_{\text{align}}$, where:
$\mathcal{L}_{\text{align}}=1-\frac{\langle R,S_{\text{align}}\rangle}{\|R\|_{2}\|S_{\text{align}}\|_{2}}$
The hyperparameters $\lambda_{1}$ and $\lambda_{2}$ balance the task loss and the regularization terms, ensuring that the model optimizes both task performance and alignment with the reference structure.
This procedure ensures that the learned steering function preserves attribute disentanglement, supports subspace coordination, and enables adaptive attribute combination. It ultimately yields a robust steering model capable of controlling multiple behavior dimensions in LLMs with minimal interference.
Algorithm 1 Multi-Subspace Representation Steering
1:Datasets $\mathcal{D}_{i}$ for attributes $i=1,\dots,n$, model $M$, token position $p$, layer $l$, $\Phi_{l,p}(h;R,W,b,m)$, $m_{\text{prior}}\in[0,1]^{\text{r}}$, regulation hyperparameters $\lambda_{1},\lambda_{2}$, label $y$
2:Steering model $\Phi_{l,p}(h;R,W,b,m)$
3:Activation-wise Preparation
4:for each attribute $i$ do
5:  $\tau_{i}=\frac{1}{|\mathcal{D}_{i}|}\sum_{j}h_{i,j}$
6:end for
7:Combine activations: $\tau_{c}=[\tau_{1},\dots,\tau_{n}]$
8:Shared Subspace Extraction
[…]
11:Private Subspace Extraction
12:for each attribute $i$ do
13:  $\tau_{i}=U_{i}\Sigma_{i}V_{i}^{\top}$
14:  $H_{\text{res}}^{(i)}=\tau_{i}-R_{\text{shared}}^{\top}(R_{\text{shared}}\tau_{i})$
15:  $H_{\text{res}}^{(i)}=U^{(i)}S^{(i)}V^{(i)\top}$
16:  $B_{i}=U_{1:r_{i}}^{(i)\,\top}$ $\triangleright$ Top $r_{i}$ directions
17:end for
18:Construct Aligned Subspace
[…]
20:Optimize Representation Matrix
21:Initialize $R$, $W$, $b$, subspace mask $m(h)$
22:for each training step do
23:  $\Phi_{l,p}(h)=h+R^{\top}\cdot\text{diag}(m(h))\cdot(Wh+b-Rh)$
24:  $\mathcal{L}_{\text{task}}=\text{CrossEntropy}\left(\text{Softmax}\left(\Phi_{l,p}(h)\right),y\right)$
25:  $\mathcal{L}_{\text{align}}=1-\frac{\langle R,S_{\text{align}}\rangle}{\|R\|_{2}\|S_{\text{align}}\|_{2}}$
26:  $\mathcal{L}=\mathcal{L}_{\text{task}}+\lambda_{1}\|m(h)-m_{\text{prior}}\|_{2}^{2}+\lambda_{2}\mathcal{L}_{\text{align}}$
27:  Update parameters $R$, $W$, $b$, $m$
28:end for
29:return $\Phi_{l,p}(h;R,W,b,m)$
@@ Appendix B Implementation Details
A.2 Datasets and Metircs We provide detailed descriptions of the datasets and evaluation metrics used in our experiments to assess multi-attribute steering performance.
@@
B.1 Performance on MMLU Benchmark
This section provides a comprehensive evaluation of MSRS on the Massive Multitask Language Understanding (MMLU) benchmark, assessing its ability to maintain and enhance general language capabilities across diverse domains. The MMLU benchmark, comprising 57 tasks across 17 subjects, serves as a rigorous testbed for evaluating model robustness beyond attribute-specific steering. We compare MSRS against baseline ITI, MTL-LoRA, and CAA on three base models: LLaMA3-8B-Instruct, Qwen2-7B-Instruct, and Mistral-7b-v0.3. Table 6 summarizes the per-subject performance of MSRS and baselines on the MMLU benchmark. The results reveal distinct patterns in how each method impacts general capabilities, with MSRS demonstrating superior consistency and enhancement over baselines.
MSRS enhances or preserves performance across subjects. On LLaMA3-8B-Instruct, MSRS achieves the highest accuracy in 14 out of 17 subjects, with significant gains in STEM fields such as mathematics (0.567 vs. 0.430 for the base model), physics (0.642 vs. 0.533), and chemistry (0.578 vs. 0.502). These improvements underscore MSRS’s ability to bolster reasoning and knowledge retention in technically demanding domains. In humanities and social sciences, such as history (0.809 vs. 0.778) and psychology (0.814 vs. 0.764), MSRS also outperforms the base model, indicating broad applicability. For Qwen2-7B-Instruct, MSRS matches or exceeds the base model’s performance in most subjects, with notable improvements in geography (0.889 vs. 0.874) and psychology (0.814 vs. 0.803). These gains, though modest, highlight MSRS’s stability across high-performing base models, preserving their strong initial capabilities while enabling targeted enhancements. On Mistral-7b-v0.3, MSRS delivers substantial uplifts in subjects like psychology (0.790 vs. 0.737) and law (0.532 vs. 0.503), despite the base model’s lower baseline performance. However, in mathematics (0.365 vs. 0.387), MSRS slightly underperforms, suggesting potential limitations in enhancing weaker base models in certain domains.
Baseline methods reveal trade-offs. ITI consistently degrades performance across most subjects due to its focus on truthfulness. On LLaMA3-8B-Instruct, ITI drops accuracy in mathematics to 0.371 (vs. 0.430) and computer science to 0.476 (vs. 0.629), reflecting a significant loss in general reasoning and knowledge. Similar declines are observed across all models, confirming ITI’s unsuitability for preserving broad capabilities. MTL-LoRA also exhibits reduced performance, particularly in STEM subjects. For Mistral-7b-v0.3, accuracy in mathematics falls to 0.318 (vs. 0.387) and chemistry to 0.386 (vs. 0.492). On LLaMA3-8B-Instruct, declines are evident in economics (0.547 vs. 0.675) and engineering (0.503 vs. 0.641). These results suggest that MTL-LoRA’s multi-task fine-tuning overfits to specific tasks, compromising the model’s general knowledge base. CAA performs closer to the base models but rarely surpasses them. On Qwen2-7B-Instruct, CAA achieves comparable scores (e.g., 0.567 in mathematics, 0.854 in business) but trails MSRS in geography (0.884 vs. 0.889) and psychology (0.813 vs. 0.814). On Mistral-7b-v0.3, CAA maintains baseline levels (e.g., 0.774 in culture) without the consistent improvements seen in MSRS, indicating limited generalization beyond attribute steering.
The experimental results affirm that MSRS excels in maintaining and often enhancing general NLP capabilities across the MMLU benchmark, outperforming baseline methods in both consistency and performance. Unlike ITI and MTL-LoRA, which sacrifice broad knowledge for attribute-specific gains, and CAA, which offers limited improvement, MSRS achieves a better balance between targeted steering and general understanding.
B.2 Performance on GLUE Benchmark
@@
MSRS outperforms baselines across GLUE tasks. For LLaMA3-8B-Instruct, MSRS achieves an average score of 0.7748, surpassing the base model (0.7257), CAA (0.7384), and ReFT (0.7569). It excels particularly in SST-2 (0.9799) and QNLI (0.8097), highlighting its strengths in sentiment classification and reasoning tasks. Additionally, MSRS improves RTE (0.6912) compared to CAA (0.6701) and ReFT (0.6577). On Qwen2-7B-Instruct, MSRS records an average score of 0.8322, slightly better than ReFT (0.8300) and notably higher than CAA (0.7701). It achieves top performance in QNLI (0.8675) and QQP (0.8672), demonstrating consistency across tasks, even though it slightly trails ReFT in SST-2 (0.8850 vs. 0.8750). With Mistral-7B-v0.3, MSRS attains an average score of 0.7066, outperforming the base model (0.6808), CAA (0.6551), and ReFT (0.6927). It shows significant gains in CoLA (0.8208) and RTE (0.6153), where baselines struggle (e.g., CAA’s RTE: 0.3225).
Baseline methods show variability. CAA tends to improve specific tasks but lacks generalization. For instance, on LLaMA3-8B-Instruct, it boosts SST-2 (0.9641) but drops in RTE (0.6701). On Mistral-7B-v0.3, CAA severely underperforms in QNLI (0.3225) and RTE (0.3225). ReFT achieves competitive results in some areas but is inconsistent. On Qwen2-7B-Instruct, it excels in QNLI (0.8637) but lags in SST-2 (0.8750). For Mistral-7B-v0.3, ReFT improves CoLA (0.8452) yet struggles with QQP (0.6398).
The GLUE benchmark results underscore the effectiveness of MSRS in enhancing language understanding across a range of tasks. MSRS consistently achieves the highest average scores across all three models, LLaMA3-8B-Instruct (0.7748), Qwen2-7B-Instruct (0.8322), and Mistral-7B-v0.3, outperforming both CAA and ReFT. Its shared subspace mechanism enables MSRS to generalize effectively, balancing task-specific improvements with broad linguistic competence.
@@ Appendix C Adaptive Subspace Selecting
Limitations of same space training. When all attributes are trained in a shared subspace, performance suffers due to interference between conflicting attribute objectives. For example, on LLaMA3-8B-Instruct, Same Space achieves an MC1 score of 29.58 on TruthfulQA, compared to 32.52 for $\text{MSRS}_{\text{Attribute}}$ and 33.50 for $\text{MSRS}_{\text{Rank}}$. Similarly, on Qwen2-7B-Instruct, it records a Sorry-bench score of 0.422, lagging behind $\text{MSRS}_{\text{Attribute}}$ at 0.446. This suggests that a single subspace struggles to accommodate diverse steering directions, resulting in suboptimal optimization.
Advantages of adaptive subspace selection. The MSRS variants address this interference by decoupling subspaces and applying adaptive weighting through the mask network $m(h)$. $\text{MSRS}_{\text{Attribute}}$: By grouping low-rank dimensions into attribute-specific subspaces, this strategy balances parameter sharing and specialization. On LLaMA3-8B-Instruct, it improves MC1 to 32.52 and BLEU to 52.57, while on Qwen2-7B-Instruct, it achieves an MC1 of 34.72 and BLEURT of 74.90, consistently surpassing Same Space. $\text{MSRS}_{\text{Rank}}$: This finer-grained approach weights each dimension individually, excelling on LLaMA3-8B-Instruct with an MC1 of 33.50 and MC2 of 52.74, indicating superior truthfulness. However, on Qwen2-7B-Instruct, it underperforms $\text{MSRS}_{\text{Attribute}}$ (MC1: 26.41 vs. 34.72), suggesting that excessive granularity may not always be beneficial.
Impact of subspace granularity. The optimal granularity varies by model. For LLaMA3-8B-Instruct, $\text{MSRS}_{\text{Rank}}$’s dimension-level control yields slight improvements over $\text{MSRS}_{\text{Attribute}}$, particularly in TruthfulQA metrics. In contrast, $\text{MSRS}_{\text{Attribute}}$ outperforms $\text{MSRS}_{\text{Rank}}$ on Qwen2-7B-Instruct and Mistral-7B-v0.3, suggesting that coarser subspace grouping aligns better with certain model architectures.
@@ Appendix D Intervention Position Experiments
LLaMA2-7B: The Important Token strategy yields notable improvements over Last Token, with MC1 increasing from 26.41 to 29.10, BLEU from 48.66 to 49.88, and BBQ accuracy from 0.631 to 0.644. On HelpSteer, it enhances Helpfulness (2.70 to 3.12), Coherence (2.68 to 3.06), and Verbosity (2.73 to 2.47), reflecting more guided and informative outputs.
LLaMA3-8B-Instruct: The Important Token approach boosts MC2 from 52.74 to 56.32 and rouge1 from 56.48 to 58.22, alongside a BBQ accuracy increase from 0.6457 to 0.6553. HelpSteer Verbosity improves from 2.24 to 1.99, though Helpfulness and Coherence remain stable, suggesting robust baseline performance.
Qwen2-7B-Instruct: Gains are observed in MC1 (34.72 to 36.12), BLEU (51.10 to 52.17), and BBQ accuracy (0.6421 to 0.6572). HelpSteer Verbosity rises from 2.17 to 2.06, with minimal changes in Helpfulness and Coherence, indicating consistent but moderate enhancements.
The experimental results validate the effectiveness of the Dynamic Intervention Position Selection mechanism. By targeting the most semantically relevant tokens, the Important Token strategy consistently outperforms the Last Token baseline across diverse datasets and models. These improvements are evident in enhanced truthfulness, reduced bias, and higher-quality generations, alongside greater interpretability of model behavior.
@@ Appendix E Steering Layer Selection
Performance Trends Across Layers Lower Layers (e.g., Layer 3): Interventions at lower layers exhibit limited steering capability. For example, LLaMA3-8B-Instruct at Layer 3 achieves an MC1 score of 29.83 and BBQ accuracy of 0.631, underperforming compared to higher layers. This can be attributed to the early layers’ focus on syntactic rather than semantic representations, limiting their effectiveness for attribute control.
Mid-to-Upper Layers (e.g., Layer 15): Optimal performance is consistently observed at mid-to-upper layers, with Layer 15 standing out across all models. For LLaMA3-8B-Instruct, Layer 15 yields an MC1 of 33.50, BLEU of 52.32, and HelpSteer Helpfulness of 3.88. Similarly, Qwen2-7B-Instruct at Layer 15 achieves an MC1 of 34.72, BLEURT of 74.90, and BBQ accuracy of 0.642. These results suggest that mid-layers strike an effective balance between semantic abstraction and model generalization.
Deeper Layers (e.g., Layer 27): Performance degrades at deeper layers, likely due to overfitting or overly specialized representations. For instance, Mistral-7B-v0.3 at Layer 27 records an MC1 of 24.83 and BLEU of 45.76, indicating a reduced capacity to generalize effectively when interventions occur late in the transformer stack.
Layer Selection via Grid Search To determine the optimal intervention layer for multi-attribute subspace training, we employ a grid search over held-out validation splits. This method systematically evaluates performance across layers and attributes, identifying Layer 15 as the most effective choice for balancing trade-offs. This targeted selection is adopted in all subsequent experiments to ensure that steering interventions maximize the utility of the learned multi-subspace representations.

### v4

@@ 6 Conclusion
We present Multi-Subspace Representation Steering (MSRS), a principled and effective framework for multi-attribute behavior control in LLMs. MSRS addresses key limitations of prior methods by introducing attribute-specific subspaces with SVD-guided dimensionality, a shared subspace to capture cross-attribute correlations for precise intervention. By integrating these components, MSRS mitigates steering conflicts and improves controllability across diverse attributes. Extensive experiments demonstrate that MSRS consistently outperforms existing approaches in both task-specific and general-purpose settings, offering a scalable and robust solution for reliable and aligned language generation.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Implementation Details
@@
A.1 Datasets and Metrics We provide detailed descriptions of the datasets and evaluation metrics used in our experiments to assess multi-attribute steering performance.
@@
SafetyBench
(Zhang et al., 2023): Measures performance on multiple safety dimensions including offensiveness, unfairness, ethics and morality, and privacy and property. This benchmark evaluates the model’s ability to generate safe and responsible responses, covering a broad range of sensitive topics. The evaluation involves assessing the degree of harmfulness in responses, and how well the model adheres to ethical and privacy considerations.
GSM8k
(Cobbe et al., 2021): Evaluates reasoning ability by testing the model on a variety of mathematical word problems. The GSM8k benchmark is designed to assess the model’s ability to perform multi-step reasoning and inference, as well as its ability to handle mathematical tasks at a high level of accuracy.
Human-Style-Answers
: A benchmark designed to evaluate the style of human-like responses generated by AI models. It focuses on generating natural, contextually appropriate, and coherent responses. The evaluation includes assessing the fluency, tone, and style of the model’s responses, ensuring they are aligned with typical human responses. Details of the benchmark can be found at https://huggingface.co/datasets/innova-ai/Human-Style-Answers.
@@
Table 6: Model Performance for Different Numbers of Attribute Settings.
r (n) truthfulqa ($\uparrow$) bbq ($\uparrow$) alpaca ($\uparrow$) refusal ($\uparrow$) helpful ($\uparrow$) coherence ($\uparrow$) verbosity ($\downarrow$) SafetyBench ($\uparrow$) gsm8k ($\uparrow$) style ($\uparrow$)
0 36.55% 0.6081 0.12 0.4911 3.76 3.41 2.33 62.13 70.32 75.35
4 (n=1) 43.54% 0.6476 0.35 0.5313 3.89 3.91 2.21 64.79 75.21 77.82
8 (n=2) 43.12% 0.6457 0.36 0.5295 3.89 3.96 2.24 64.84 74.46 77.81
12 (n=3) 39.57% 0.6471 0.37 0.5321 3.83 3.93 2.22 64.43 73.71 78.02
15 (n=4) 40.52% 0.6301 0.33 0.5301 3.82 3.85 2.22 63.87 73.07 77.65
17 (n=5) 38.74% 0.6246 0.26 0.5179 3.82 3.85 2.23 63.38 72.35 77.41
18 (n=6) 38.52% 0.6237 0.23 0.5172 3.82 3.73 2.29 66.26 72.18 77.21
20 (n=7) 38.47% 0.6233 0.14 0.5105 3.82 3.72 2.32 62.51 71.97 76.68
21 (n=8) 37.87% 0.6197 0.12 0.5087 3.79 3.60 2.35 62.30 71.72 76.73
23 (n=9) 37.68% 0.6102 0.11 0.5091 3.79 3.58 2.36 62.28 71.73 76.81
24 (n=10) 36.88% 0.5920 0.08 0.5031 3.79 3.47 2.37 62.12 70.05 76.69
A.1 Performance on MMLU Benchmark
This section provides a comprehensive evaluation of MSRS on the Massive Multitask Language Understanding (MMLU) benchmark, assessing its ability to maintain and enhance general language capabilities across diverse domains. The MMLU benchmark, comprising 57 tasks across 17 subjects, serves as a rigorous testbed for evaluating model robustness beyond attribute-specific steering. We compare MSRS against baseline ITI and CAA on three base models: LLaMA3-8B-Instruct, Qwen2-7B-Instruct, and Mistral-7b-v0.3. Table 5 summarizes the per-subject performance of MSRS and baselines on the MMLU benchmark. The results reveal distinct patterns in how each method impacts general capabilities, with MSRS demonstrating superior consistency and enhancement over baselines.
A.2 Performance on GLUE Benchmark
@@
Appendix B Computational Complexity and Runtime Analysis
All experiments are conducted on a single NVIDIA V100 32 GiB GPU. No specialized optimizations (e.g., randomized SVD or CPU offloading) are used. We analyze the computational cost of the subspace extraction phases:
B.1 Offline Subspace Extraction
This phase is performed once per model and attribute combination.
Shared Subspace Extraction. We first forward each balanced dataset $\mathcal{D}_{i}$ ($|\mathcal{D}_{i}|\leq 1000$) through the frozen LLM and compute the mean activation vector $\tau_{i}\in\mathbb{R}^{d}$ ($d=4096$) from the last few tokens. These vectors are concatenated into:
$\tau_{c}=[\tau_{1}\,|\,\tau_{2}\,|\,\dots\,|\,\tau_{n}]\in\mathbb{R}^{d\times n}.$
A single full SVD is performed on $\tau_{c}$ with complexity $\mathcal{O}(dn^{2})$.
Iterative Attribute-Specific Subspace Extraction. For each attribute $i$, we form the per-sample activation matrix
$H_{i}^{(i)}=[h_{i,1}^{l},\dots,h_{i,|\mathcal{D}_{i}|}^{l}]\in\mathbb{R}^{d\times|\mathcal{D}_{i}|},\quad|\mathcal{D}_{i}|\leq 1000.$
To ensure orthogonality among all learned subspaces, we iteratively subtract the projections of the shared subspace and all previously extracted attribute subspaces. The residual is computed as:
$H_{\text{res}}^{(i)}=H_{i}^{(i)}-B_{\text{shared}}^{\top}B_{\text{shared}}H_{i}^{(i)}-\sum_{k=1}^{i-1}B_{k}^{\top}B_{k}H_{i}^{(i)},$ (6)
where $B_{\text{shared}}$ and $B_{k}$ are the orthonormal bases for the shared and prior attribute subspaces ($k<i$), respectively. We then perform full SVD on each $H_{\text{res}}^{(i)}$ independently. Theoretical complexity per attribute: $\mathcal{O}(d\cdot|\mathcal{D}_{i}|^{2})\approx\mathcal{O}(4{,}096\times 10^{6})\approx 4\times 10^{9}$ FLOPs in the worst case.
Measured Wall-Clock Time
Table 8: Offline subspace extraction time on Llama-3-8B for $n$ attributes and NN samples per attribute.
Phase Matrix size Time per run (seconds) Total Time (seconds)
Activation extraction - 0.29 (per sample) $N\times n\times 0.29$
Shared subspace $4096\times n$ - 0.18
Per-attribute subspace $4096\times\leq\!1000$ each 3.42 (per attribute) $3.42\times n$
Full offline phase $0.18+(3.42\times n)+(N\times n\times 0.29)$
Realistic estimate for the setting ($n=3$, with $N=1000$ samples per attribute):
$\text{Total time}\approx 0.18+(3.42\times 3)+(1000\times 3\times 0.29)\approx 0.18+10.26+870=880.44\text{ seconds}$
This cost is incurred only once per model and attribute set. The vast majority of this time is spent on forward passes to collect activations. The resulting subspace bases $B_{\text{shared}}$ and $\{B_{i}\}_{i=1}^{n}$ are saved and directly loaded by all subsequent runs, introducing zero additional overhead during online training or inference.
B.2 Online Training Cost
Following the subspace extraction, we proceed to the online training phase where the steering parameters (e.g., the mask network $m_{\theta}$ and projection matrices) are optimized.
Experimental Setup. We train MSRS using the same hardware (single V100 32GB) with the following hyperparameters: rank $r=8$, batch size $B=4$, and a dataset size of $N=500$ samples per attribute. The model is trained for $E=10$ epochs.
Complexity & Overhead. The computational graph of MSRS is nearly identical to standard ReFT (Low-rank Linear Subspace ReFT), as the additional mask prediction network is a lightweight MLP ($d\to n$) with negligible FLOPs compared to the Transformer backbone’s forward/backward pass.
Measured Training Time. Table 9 summarizes the training throughput. With gradient checkpointing enabled, the average time per training step is approximately 0.45 seconds.
Table 9: Online training cost comparison on Llama-3-8B ($n=3$ attributes, $N=500$ per attribute).
Method Params Time per Step Total Steps Total Training Time
($B=4$) ($E=10$) (Seconds) (GPU Hours)
LoReFT (Baseline) $d\times r$ 0.42s 3,750 $\approx 1,575$s $\approx 0.44$h
MSRS (Ours) $d\times r+\text{Mask}$ 0.45s 3,750 $\approx 1,688$s $\approx\mathbf{0.47}$h
For this setting ($n=3$, total samples $1,500$), the total training budget is:
$\text{Cost}=E\times\frac{N_{\text{total}}}{B}\times T_{\text{step}}\approx 10\times\frac{1500}{4}\times 0.45s\approx 1688\text{ seconds}\approx\mathbf{0.47\text{ GPU Hours}}.$
This demonstrates that MSRS requires less than 0.5 GPU hours to converge, making it highly efficient for rapid experimentation and deployment.
@@ Appendix E Steering Layer Selection
Layer Selection via Grid Search To determine the optimal intervention layer for multi-attribute subspace training, we employ a grid search over held-out validation splits. This method systematically evaluates performance across layers and attributes, identifying the most effective layer for balancing trade-offs. This targeted selection is adopted in all subsequent experiments to ensure that steering interventions maximize the utility of the learned multi-subspace representations.